### v1

# Differential calculus on Hopf–Galois extension via the Đurđević braiding

Abstract
We introduce a class of first–order differential calculus on principal comodule algebras generated by the Đurđević braiding $\sigma$ and a chosen vertical ideal. Starting from the universal calculus and a strong connection, we construct $\sigma$–generated calculus and prove their existence for arbitrary principal comodule algebras. We show that, in this setting, connection 11–forms and vertical maps descend to the quotient calculus and are compatible with the induced braided symmetry. We also compare this framework with Đurđević’s complete differential calculus.
1 Introduction
Noncommutative differential geometry provides a natural framework for studying geometric structures on quantum spaces arising from Hopf algebras and quantum groups. A central class of examples is furnished by quantum principal bundles, which are algebraic counterparts of classical principal fiber bundles and are modelled by Hopf–Galois extensions. Since the foundational work of Brzeziński–Majid [4], Hajac [12], Brzeziński–Hajac [6] such extensions equipped with strong connections have been recognized as appropriate noncommutative analogues of principal bundles, supporting gauge theory, Chern–Connes pairings, and index-theoretic constructions [7, 13].
In this framework, differential calculi play a fundamental role. They encode infinitesimal geometric data and provide the setting for connections, curvature, and gauge transformations. Given an algebra $A$, every first–order differential calculus can be realized as a quotient of the universal calculus $\Omega^{1}_{\mathrm{u}}(A)$ by a suitable AA–subbimodule [15]. For Hopf–Galois extensions, additional covariance and compatibility conditions are required in order to reflect the underlying bundle structure.
An important contribution to this subject is due to Đurđević, who introduced a braided symmetry on quantum principal bundles in terms of a canonical braiding on the balanced tensor product $A\otimes_{A^{\operatorname{co}H}}A$ [11]. This braiding, now known as the Đurđević braiding, governs the interaction between horizontal and vertical directions and plays a central role in the description of gauge transformations and braided symmetries [1, 2].
Building on this idea, Đurđević introduced the notion of a complete differential calculus on a quantum principal bundle, consisting of a graded differential algebra equipped with an extended coaction and satisfying strong compatibility conditions [9, 10]. In this setting, the noncommutative Atiyah sequence is exact in all degrees and admits a canonical horizontal–vertical decomposition. More recently, this framework has been further analysed and refined in [8], where structural properties and applications were investigated in detail.
While complete differential calculi provide a powerful and geometrically rich theory, their construction requires substantial higher–order data and strong regularity assumptions. In many situations, however, one is primarily interested in first–order structures and in the behaviour of connection forms and vertical maps. This motivates the search for weaker frameworks that retain the essential braided and descent properties while avoiding the full complexity of complete calculi.
The purpose of this paper is to introduce and study a class of first–order differential calculi on principal comodule algebras that are generated by the Đurđević braiding and a chosen vertical ideal. More precisely, given a principal right HH–comodule algebra $A$ with a strong connection and a right ideal $I_{H}\subset H^{+}=\ker\varepsilon$, we consider differential calculi whose defining relations are generated by the image of the associated universal connection 11–form and closed under the action of the Đurđević braiding. We refer to such calculi as $\sigma$–generated, where $\sigma$ denotes here the Đurđević braiding.
The notion of $\sigma$–generation isolates a minimal first–order structure underlying the vertical geometry of Hopf–Galois extensions. It ensures that connection 11–forms and vertical maps descend from the universal calculus to the quotient calculus and are compatible with the intrinsic braided symmetry (parallel treatment can be found in [8], Lemma 4.7). At the same time, no horizontal–vertical decomposition or higher–order compatibility is imposed.
The main results of this paper may be summarized as follows. First, we show that on any principal comodule algebra, and for any fixed right ideal $I_{H}\subset H^{+}$, there exists a $\sigma$–generated differential calculus. This provides a systematic method for constructing braided-compatible first–order calculi from strong connections. Second, we analyse several necessary conditions for $\sigma$–generation and identify obstructions that rule out this property for general covariant calculi. Third, we establish that for $\sigma$–generated calculi, universal connection 11–forms and vertical maps descend to the quotient calculus and are compatible with the Đurđević braiding. We also clarify the precise relationship between $\sigma$–generated calculi and complete differential calculi in the sense of Đurđević.
Our approach differs from the complete calculus framework in that we work exclusively at the level of first–order forms and formulate all conditions in terms of the universal calculus, strong connections, and the induced braiding. This allows us to treat arbitrary principal comodule algebras and to describe explicitly the defining relations of the resulting calculi. From this perspective, $\sigma$–generated calculi may be viewed as first–order approximations of the complete calculus formalism, retaining its essential braided and descent properties while remaining applicable in greater generality.
The paper is organized as follows. In Section 2, we recall basic notions concerning differential calculi, Hopf–Galois extensions, principal comodule algebras, and the Đurđević braiding. In Section 3, we introduce $\sigma$–generated differential calculi and establish their basic properties. We compare our framework with the theory of complete differential calculi. Section 4 is devoted to necessary conditions and obstructions to $\sigma$–generation. Finally in Section 5, we present an example of quantum Hopf fibration over standard Podleś sphere and an associated $\sigma$-generated differential calculus.
2 Preliminaries
We fix a field $k$. Throughout this paper, all spaces are $k$-vector spaces and all maps are $k$-linear unless otherwise stated. The tensor product of $k$-spaces is denoted by $\otimes$, and the balanced tensor product over an algebra $B$ is denoted by $\otimes_{B}$. We denote $H$ to be a Hopf algebra with coproduct $\Delta$, counit $\varepsilon$ and a bijective antipode $S$. An algebra $A$ is always assumed to be associative and unital. We use Sweedler notation for coproducts and coactions, suppressing summation symbols. In this section we follow the literature from [3, 5, 11].
2.1 Universal and first–order differential calculi
The universal first–order differential calculus over an algebra $A$ is defined as
$\Omega^{1}_{\mathrm{u}}(A):=\ker(m:A\otimes A\to A),$
with universal differential $\mathrm{d}_{\mathrm{u}}:A\to\Omega^{1}_{\mathrm{u}}(A)$ given by
$\mathrm{d}_{\mathrm{u}}(a)=1\otimes a-a\otimes 1.$
A first–order differential calculus over $A$ is a pair $(\Omega^{1}(A),\mathrm{d})$ such that $\Omega^{1}(A)$ is an AA–bimodule and $\mathrm{d}:A\to\Omega^{1}(A)$ satisfies the Leibniz rule and $\Omega^{1}(A)=\mathrm{span}\{a\,\mathrm{d}b:a,b\in A\}$. Every first–order differential calculus arises as a quotient
$\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A},$
for some AA–subbimodule $N_{A}\subset\Omega^{1}_{\mathrm{u}}(A)$.
2.2 Principal comodule algebra
Let $H$ be a Hopf algebra and $A$ be a right HH–comodule algebra with coaction $\delta:A\to A\otimes H$. The subalgebra of coinvariants is
$B:=A^{\operatorname{co}(H)}=\{\,b\in A\mid\delta(b)=b\otimes 1\,\}.$
The extension $B\subset A$ is called a Hopf–Galois extension if the canonical map
$\mathrm{can}:A\otimes_{B}A\longrightarrow A\otimes H,\qquad a\otimes_{B}a^{\prime}\longmapsto a\,\delta(a^{\prime})=aa^{\prime}_{(0)}\otimes a^{\prime}_{(1)},$
is bijective. In this case one defines the translation map
$\tau:H\longrightarrow A\otimes_{B}A,\qquad\tau(h)=\mathrm{can}^{-1}(1\otimes h),\qquad\tau(h)=h^{\langle 1\rangle}\otimes_{B}h^{\langle 2\rangle}.$
Definition 2.1.
A right HH–comodule algebra $A$ is called a principal comodule algebra if:
  1. 1.
    $B=A^{\operatorname{co}(H)}\subset A$ is a Hopf–Galois extension;
  2. 2.
    the multiplication map $\mu:B\otimes A\to A$ admits a left BB–linear, right HH–colinear splitting $s:A\rightarrow B\otimes A$ such that $\mu\circ s=\mathrm{id}_{A}$.
2.3 Vertical map and strong connection
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension.
Definition 2.2.
The vertical map on a universal calculus is a linear map
$\mathrm{ver}_{\mathrm{u}}:\Omega^{1}_{\mathrm{u}}(A)\longrightarrow A\otimes H^{+}$
defined by
$\mathrm{ver}_{\mathrm{u}}(a\otimes a^{\prime})=aa^{\prime}_{(0)}\otimes a^{\prime}_{(1)},$
where $H^{+}=\ker\varepsilon$.
Definition 2.3.
A strong connection on a Hopf-Galois extension $B=A^{\mathrm{co}(H)}\subset A$ is a linear map
$\ell:H\longrightarrow A\otimes A$
satisfying
  1. 1.
    $\ell(1_{H})=1_{A}\otimes 1_{A}$;
  2. 2.
    $(\mathrm{id}_{A}\otimes\delta)\circ\ell=(\ell\otimes\mathrm{id}_{H})\circ\Delta$;
  3. 3.
    $\widetilde{\mathrm{can}}\circ\ell=1_{A}\otimes\mathrm{id}_{H}$, where $\widetilde{\mathrm{can}}:A\otimes A\rightarrow A\otimes H$ be the lift of $\mathrm{can}$ to $A\otimes A$.
Definition 2.4.
Let $\ell:H\to A\otimes A$ be a strong connection. The associated connection 11–form is the linear map
$\omega:H^{+}\longrightarrow\Omega^{1}_{\mathrm{u}}(A),\qquad\omega(h^{+})=\ell(h)-\varepsilon(h)\,1_{A}\otimes 1_{A},$
where $H^{+}=\ker\varepsilon$, and $h^{+}:=h-\varepsilon(h)1_{H}$. It satisfies the following properties:
  1. 1.
    $\mathrm{ver}_{\mathrm{u}}\circ\omega=1_{A}\otimes\mathrm{id}_{H^{+}}$;
  2. 2.
    $\delta_{\Omega^{1}_{\mathrm{u}}(A)}\circ\omega=(\omega\otimes\mathrm{id}_{H})\circ\mathrm{Ad}_{R}$, where $\delta_{\Omega^{1}_{\mathrm{u}}(A)}$ denotes the right HH–coaction on $\Omega^{1}_{\mathrm{u}}(A)$ and $\mathrm{Ad}_{R}$ denotes the right adjoint coaction of $H$ on $H^{+}$.
By [[6], Theorem 2.5], every principal comodule algebra admits a strong connection.
2.4 Đurđević braiding
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension. Using the translation map $\tau$, one defines the Đurđević braiding
$\sigma:A\otimes_{B}A\longrightarrow A\otimes_{B}A$
by
$\sigma(a\otimes_{B}a^{\prime})=a_{(0)}a^{\prime}\,\tau(a_{(1)})=a_{(0)}a^{\prime}(a_{(1)})^{\langle 1\rangle}\otimes_{B}(a_{(1)})^{\langle 2\rangle}.$
Proposition 2.5 ([11], Proposition 2.1).
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf-Galois extension. Then the following hold true:
  1. 1.
    The map $\sigma:A\otimes_{B}A\rightarrow A\otimes_{B}A$ is an isomorphism. Its inverse is
    $\sigma^{-1}(a\otimes_{B}a^{\prime})=\tau(S^{-1}(a^{\prime}_{(1)}))aa^{\prime}_{(0)}=S^{-1}(a^{\prime}_{(1)})^{\langle 1\rangle}\otimes_{B}S^{-1}(a^{\prime}_{(1)})^{\langle 2\rangle}aa^{\prime}_{(0)}$
  2. 2.
    $\sigma$ satisfies the braid relation:
    $(\sigma\otimes_{B}\mathrm{id})\circ(\mathrm{id}\otimes_{B}\sigma)\circ(\sigma\otimes_{B}\mathrm{id})=(\mathrm{id}\otimes_{B}\sigma)\circ(\sigma\otimes_{B}\mathrm{id})\circ(\mathrm{id}\otimes_{B}\sigma)$
    on $A\otimes_{B}A\otimes_{B}A$;
  3. 3.
    AA is $\sigma$-commutative, in other words,
    $m_{A}\circ\sigma=m_{A}$
    where $m_{A}$ denotes the multiplication $A\otimes_{B}A\rightarrow A$;
  4. 4.
    The product map $m_{A}$ is compatible with braiding $\sigma$, that is
    1. (a)
      $\sigma\circ(m_{A}\otimes_{B}\mathrm{id})=(\mathrm{id}\otimes_{B}m_{A})\circ(\sigma\otimes_{B}\mathrm{id})\circ(\mathrm{id}\otimes_{B}\sigma)$
    2. (b)
      $\sigma\circ(\mathrm{id}\otimes_{B}m_{A})=(m_{A}\otimes_{B}\mathrm{id})\circ(\mathrm{id}\otimes_{B}\sigma)\circ(\sigma\otimes_{B}\mathrm{id})$
3 Differential calculus generated by the Đurđević braiding
Let $H$ be a Hopf algebra with bijective antipode, and let $A$ be a right $H$-comodule algebra with coaction $\delta:A\to A\otimes H$. Let $B=A^{\mathrm{co}H}\subset A$ is a Hopf–Galois extension. Denote by
$\pi:A\otimes A\longrightarrow A\otimes_{B}A$
the canonical quotient map, and define the balanced image of $N_{A}$ by
$N_{A}^{\mathrm{bal}}\;:=\;\pi(N_{A})\;=\;\frac{N_{A}+A\Omega^{1}_{\mathrm{u}}(B)A}{A\Omega^{1}_{\mathrm{u}}(B)A}\;\subset\;A\otimes_{B}A.$
Lemma 3.1.
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $I_{H}\subset H^{+}$ be a right ideal. Let
$\mathrm{ver}_{\mathrm{u}}:\Omega^{1}_{\mathrm{u}}(A)\longrightarrow A\otimes H^{+}$
be the vertical map,
$\pi:\Omega^{1}_{\mathrm{u}}(A)\longrightarrow A\otimes_{B}A$
the canonical projection, and
$\mathrm{can}:A\otimes_{B}A\longrightarrow A\otimes H$
the Hopf–Galois canonical map. Then
$\pi\big(\mathrm{ver}_{u}^{-1}(A\otimes I_{H})\big)\;=\;\mathrm{can}^{-1}(A\otimes I_{H})\;\subset\;A\otimes_{B}A.$
Proof.
Let $\mathrm{pr}_{H^{+}}:H\to H^{+}$ denote the canonical projection $\mathrm{pr}_{H^{+}}(h)=h-\varepsilon(h)1_{H}$. By definition of the vertical map one has
$\mathrm{ver}_{u}=(\mathrm{id}\otimes\mathrm{pr}_{H^{+}})\circ\mathrm{can}\circ\pi$
Let $x\in\Omega^{1}_{\mathrm{u}}(A)$. Since $\Omega^{1}_{\mathrm{u}}(A)=\ker(m)$, it follows that
$(\mathrm{id}\otimes\varepsilon)\big(\mathrm{can}(\pi(x))\big)=0$
Indeed, for $x=a\,\mathrm{d}_{\mathrm{u}}b$ one has $\mathrm{can}(\pi(x))=ab_{(0)}\otimes b_{(1)}-ab\otimes 1$, and applying $\mathrm{id}\otimes\varepsilon$ yields zero.
Combining (1) and (2) gives
$\mathrm{ver}_{\mathrm{u}}(x)=\mathrm{can}(\pi(x)).$
Since $I_{H}\subset H^{+}$, equality (3) implies
$\mathrm{ver}_{\mathrm{u}}(x)\in A\otimes I_{H}\quad\Longleftrightarrow\quad\mathrm{can}(\pi(x))\in A\otimes I_{H}.$
Applying $\pi$ and using the bijectivity of $\mathrm{can}$ yields
$\pi\big(\mathrm{ver}_{\mathrm{u}}^{-1}(A\otimes I_{H})\big)=\mathrm{can}^{-1}(A\otimes I_{H}),$
as claimed. ∎
Remark 3.2.
Throughout, $I_{H}\subset H^{+}$ denotes a fixed right ideal specifying vertical relations. We neither assume a differential calculus on $H$, nor adjoint coaction invariance of $I_{H}$. Such additional conditions arise only in the framework of quantum principal bundles with compatible calculi on $A$ and $H$, which we do not assume here.
Lemma 3.3.
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $I_{H}\subset H^{+}$ be a right ideal of $H$. Let
$\mathrm{ver}_{\mathrm{u}}:\Omega^{1}_{\mathrm{u}}(A)\longrightarrow A\otimes H^{+}$
be the vertical map on the universal calculus, and
$\pi:\Omega^{1}_{\mathrm{u}}(A)\longrightarrow A\otimes_{B}A$
the canonical projection. Define
$V_{I_{H}}\;:=\;\pi\big(\mathrm{ver}_{\mathrm{u}}^{-1}(A\otimes I_{H})\big)\;\subset\;A\otimes_{B}A.$
Then $V_{I_{H}}$ is invariant under the Đurđević braiding $\sigma$, that is,
$\sigma(V_{I_{H}})\subseteq V_{I_{H}}.$
Proof.
By Lemma 3.1 one has
$V_{I_{H}}=\mathrm{can}^{-1}(A\otimes I_{H})\;\subset\;A\otimes_{B}A.$
Let $y=u\otimes_{B}a\in V_{I_{H}}$ for $u,a\in A$. Then $\mathrm{can}(y)\in A\otimes I_{H}$. Since $\mathrm{can}$ is bijective, it suffices to show that
$\mathrm{can}\big(\sigma(y)\big)\in A\otimes I_{H}.$
Recall that the Đurđević braiding $\sigma:A\otimes_{B}A\to A\otimes_{B}A$ is defined by
$\sigma(u\otimes_{B}a)=u_{(0)}\,a\,u_{(1)}^{\langle 1\rangle}\;\otimes_{B}\;u_{(1)}^{\langle 2\rangle},$
The defining property of $\tau$ is
$\mathrm{can}\big(\tau(h)\big)=1\otimes h,\qquad h\in H.$
Applying $\mathrm{can}$ to $\sigma(u\otimes_{B}a)$ and using (6) yields
$\mathrm{can}\big(\sigma(u\otimes_{B}a)\big)$ $=u_{(0)}\,a\,u_{(1)}^{\langle 1\rangle}\;u_{(1)}^{\langle 2\rangle}{}_{(0)}\;\otimes\;u_{(1)}^{\langle 2\rangle}{}_{(1)}$
$=u_{(0)}\,a\otimes u_{(1)}.$
By using the left $A$-linearity of the coaction of $\delta:A\rightarrow A\otimes H$, we have
$ua_{(0)}\otimes a_{(1)}=u_{(0)}a_{(0)}\otimes u_{(1)}a_{(1)}.$
Since $\mathrm{can}(y)\in A\otimes I_{H}$, this implies $a_{(1)}\in I_{H}$. This again implies from the identity (8), $u_{(1)}a_{(1)}\in I_{H}$. As $I_{H}$ is an ideal of $H$, we further obtain $u_{(1)}a_{(1)}S(a_{(2)})\in I_{H}$, and hence, $u_{(1)}\in I_{H}$. So, from the identity (7), it follows that $\mathrm{can}(\sigma(u\otimes_{B}a))\in A\otimes I_{H}$, which proves the claim. ∎
Remark 3.4.
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $I_{H}\subset H^{+}$ be a right ideal of $H$. Let $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)$ be the connection $1$-form associated to a strong connection on $A$. Then
$\pi(\omega(I_{H}))\subset V_{I_{H}}.$
Proof.
For any $h^{+}\in I_{H}$, the defining property of a connection 1-form gives
$\mathrm{ver}_{\mathrm{u}}(\omega(h^{+}))=1\otimes h^{+}\in A\otimes I_{H}.$
Hence $\omega(h^{+})\in\mathrm{ver}_{\mathrm{u}}^{-1}(A\otimes I_{H})$, and applying the canonical projection $\pi$ yields
$\pi(\omega(h^{+}))\in V_{I_{H}}.$
Since $h^{+}\in I_{H}$ was arbitrary, the claim follows. ∎
Definition 3.5.
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $\sigma:A\otimes_{B}A\to A\otimes_{B}A$ be the Đurđević braiding. For a subspace $W\subset A\otimes_{B}A$, denote by $\langle W\rangle_{\sigma}$ the smallest $\sigma$-stable subspace of $A\otimes_{B}A$ containing $W$. Explicitly,
$\langle W\rangle_{\sigma}:=\sum_{n\geq 0}\sigma^{n}(W).$
Definition 3.6.
Let $A$ be a principal right $H$-comodule algebra, let $I_{H}\subset H^{+}$ be a fixed right ideal of $H$, and let
$\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A}$
be a right $H$-covariant first-order differential calculus. The calculus $\Omega^{1}(A)$ is called $\sigma$-generated with respect to $I_{H}$ if there exists a strong connection $\ell:H\to A\otimes A$ with associated connection $1$-form
$\omega(h^{+})=\ell(h)-\varepsilon(h)\,1_{A}\otimes 1_{A}$
such that
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma},\qquad N_{A}^{\mathrm{bal}}:=\pi(N_{A}).$
Proposition 3.7.
Let $A$ be a principal right HH–comodule algebra and let $I_{H}\subset H^{+}$ be any fixed right ideal of $H$. Then there exists a right HH–covariant first-order differential calculus
$\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A}$
on $A$ which is $\sigma$–generated with respect to $I_{H}$, i.e.
$N_{A}^{\mathrm{bal}}=\big\langle\pi(\omega(I_{H}))\big\rangle_{\sigma}\;\subset\;A\otimes_{A^{\mathrm{co}(H)}}A,$
where $\sigma$ denotes the Đurđević braiding and $\omega$ is a connection 11–form associated to a strong connection on $A$.
Proof.
Since $A$ is a principal right HH–comodule algebra, it admits a strong connection $\ell:H\to A\otimes A$. Let $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)\subset A\otimes A$ be the associated universal connection 11–form, defined by
$\omega(h^{+})=\ell(h)-\varepsilon(h)\,1_{A}\otimes 1_{A}.$
Let $\pi:A\otimes A\to A\otimes_{A^{\mathrm{co}(H)}}A$ denote the canonical projection and define
$N_{A}^{\mathrm{bal}}:=\big\langle\pi(\omega(I_{H}))\big\rangle_{\sigma},$
the smallest $\sigma$–stable subspace of $A\otimes_{A^{\operatorname{co}H}}A$ containing $\pi(\omega(I_{H}))$.
By construction, $N_{A}^{\mathrm{bal}}$ is $\sigma$–stable and consists of vertical elements. Fix a right HH–covariant AA–subbimodule $N_{A}\subset\Omega^{1}_{\mathrm{u}}(A)$ whose image under $\pi$ equals $N_{A}^{\mathrm{bal}}$. Such a lift exists since $\pi$ is a surjective AA–bimodule map.
Define
$\Omega^{1}(A):=\Omega^{1}_{\mathrm{u}}(A)/N_{A}.$
By construction, $\Omega^{1}(A)$ is a right HH–covariant first-order differential calculus on $A$, and its balanced defining relations satisfy
$N_{A}^{\mathrm{bal}}=\big\langle\pi(\omega(I_{H}))\big\rangle_{\sigma}.$
Hence $\Omega^{1}(A)$ is $\sigma$–generated with respect to $I_{H}$. ∎
Proposition 3.8.
Let $A$ be a principal right HH–comodule algebra, let $I_{H}\subset H^{+}$ be a fixed right ideal of $H$, and let
$\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A}$
be a right $H$-covariant first-order differential calculus. Assume that $\Omega^{1}(A)$ is $\sigma$-generated with respect to $I_{H}$, i.e.
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}$
for some connection $1$-form $\omega$ associated to a strong connection $\ell$ on $A$. Then the Đurđević braiding $\sigma$ induces a well-defined endomorphism
$\bar{\sigma}:(A\otimes_{B}A)/N_{A}^{\mathrm{bal}}\longrightarrow(A\otimes_{B}A)/N_{A}^{\mathrm{bal}}$
which preserves the vertical subspace
$V_{I_{H}}/N_{A}^{\mathrm{bal}},\qquad V_{I_{H}}:=\mathrm{can}^{-1}(A\otimes I_{H}).$
Proof.
Since $\Omega^{1}(A)$ is $\sigma$-generated, its balanced kernel satisfies
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}$
for a connection $1$-form $\omega$ associated to strong connection $\ell$. By Lemma 3.3 and Remark 3.4, one has
$\pi(\omega(I_{H}))\subset V_{I_{H}},$
and since $V_{I_{H}}$ is $\sigma$-stable, it follows that
$N_{A}^{\mathrm{bal}}\subset V_{I_{H}}.$
Because $N_{A}^{\mathrm{bal}}$ is $\sigma$-stable by construction, the braiding $\sigma$ descends to a well-defined endomorphism
$\bar{\sigma}:(A\otimes_{B}A)/N_{A}^{\mathrm{bal}}\longrightarrow(A\otimes_{B}A)/N_{A}^{\mathrm{bal}}.$
By Lemma 3.3, $V_{I_{H}}$ is $\sigma$-stable. Together with the inclusion (9), this implies
$\sigma(V_{I_{H}})\subset V_{I_{H}}\quad\text{and}\quad\sigma(N_{A}^{\mathrm{bal}})\subset N_{A}^{\mathrm{bal}}.$
Therefore, $\bar{\sigma}$ preserves the quotient $V_{I_{H}}/N_{A}^{\mathrm{bal}}$. ∎
Theorem 3.9.
Let $A$ be a principal right $H$-comodule algebra, let $I_{H}\subset H^{+}$ be a fixed right ideal, and let
$\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A}$
be a right $H$-covariant first-order differential calculus on $A$. Assume that $\Omega^{1}(A)$ is $\sigma$-generated with respect to $I_{H}$, i.e.
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}$
for some strong connection $\ell:H\to A\otimes A$ with associated connection $1$-form
$\omega(h^{+})=\ell(h)-\varepsilon(h)\,1\otimes 1.$
Then:
  1. 1.
    the connection $1$-form $\omega$ descends to a well-defined map
    $\bar{\omega}:H^{+}/I_{H}\longrightarrow\Omega^{1}(A);$
  2. 2.
    the vertical map descends to a surjection
    $\mathrm{ver}:\Omega^{1}(A)\longrightarrow A\otimes H^{+}/I_{H};$
Proof.
(1) Since $\Omega^{1}(A)$ is $\sigma$-generated, by definition one has
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}.$
In particular, $\pi(\omega(h^{+}))\in N_{A}^{\mathrm{bal}}$ for all $h^{+}\in I_{H}$, which implies $\omega(I_{H})\subset N_{A}$. Therefore the connection $1$-form $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)$ vanishes on $I_{H}$ modulo $N_{A}$ and hence descends uniquely to a linear map
$\bar{\omega}:H^{+}/I_{H}\longrightarrow\Omega^{1}(A),$
as claimed.
(2) The universal vertical map
$\mathrm{ver}_{\mathrm{u}}:\Omega^{1}_{\mathrm{u}}(A)\longrightarrow A\otimes H^{+}$
satisfies $\mathrm{ver}_{\mathrm{u}}(\omega(h^{+}))=1\otimes h^{+}$ for all $h^{+}\in H^{+}$. Since $\omega(I_{H})\subset N_{A}$ and $I_{H}\subset H^{+}$, it follows that
$\mathrm{ver}_{\mathrm{u}}(N_{A})\subset A\otimes I_{H}.$
Consequently, $\mathrm{ver}_{\mathrm{u}}$ factors through the quotient $\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A}$ and induces a well-defined map
$\mathrm{ver}:\Omega^{1}(A)\longrightarrow A\otimes H^{+}/I_{H}.$
Surjectivity follows from the existence of the descended connection $1$-form $\bar{\omega}$, since $\mathrm{ver}(\bar{\omega}([h^{+}]))=1\otimes[h^{+}]$ and $\mathrm{ver}$ is left $A$-linear.
This completes the proof. ∎
Remark 3.10.
The notion of $\sigma$–generated calculus is not designed to enforce exactness of the Atiyah sequence. Rather, $\sigma$–generation provides a natural and intrinsic mechanism for the descent of vertical data and for compatibility of the resulting structures with the Đurđević braiding. Exactness in the middle of the Atiyah sequence depends on additional conditions on the horizontal relations of the calculus and is independent of $\sigma$–generation.
Relation to Đurđević’s complete differential calculus
The notion of a $\sigma$–generated first–order differential calculus introduced in this work is closely related to the complete differential calculus developed by Đurđević for Hopf–Galois extensions and its recent developments [8, 9, 11]. Since both frameworks are based on the Hopf–Galois structure and the associated Đurđević braiding, it is natural to compare their scope and structural features.
Đurđević’s complete differential calculus consists of a differential graded algebra $\Omega^{\bullet}(A)$ over a Hopf–Galois extension $B=A^{\operatorname{co}(H)}\subset A$, equipped with an extension of the right $H$ coaction to all degrees and satisfying strong compatibility conditions. In particular, the existence of a strong connection induces a canonical decomposition of first–order forms into horizontal and vertical components, yielding an exact noncommutative Atiyah sequence. In this setting, exactness and the horizontal–vertical splitting are intrinsic features of the construction.
By contrast, a $\sigma$–generated calculus is defined purely at the level of first–order forms. Starting from the universal calculus $\Omega^{1}_{\mathrm{u}}(A)$ and a strong connection with associated connection 11–form $\omega$, one imposes relations generated by $\omega(I_{H})$ and closed under the action of the Đurđević braiding $\sigma$. This guarantees the descent of connection 11–forms and vertical maps and ensures compatibility with the braided structure, but no horizontal–vertical decomposition of forms is assumed or constructed.
This distinction accounts for a fundamental difference between the two approaches. While $\sigma$–generation guarantees descent phenomena and braided compatibility, it does not, in general, imply exactness of the noncommutative Atiyah sequence. Exactness requires a mechanism for decomposing first–order forms into horizontal and vertical components, a feature that is built into Đurđević’s complete calculus but deliberately not imposed in the $\sigma$–generated setting. The possible failure of exactness therefore reflects the greater flexibility of $\sigma$–generated calculi rather than a deficiency of the construction.
Complete differential calculi are constructed using bicovariant calculi on the structure Hopf algebra and therefore involve right ideal coideals $I_{H}\subset H^{+}$ controlling the vertical component. In this framework, the defining relations of $\Omega^{1}(A)$ are tightly linked to the vertical geometry and the associated braided structure. In all known examples arising from complete calculi, the induced first–order calculus satisfies the $\sigma$–generation property. However, to the best of our knowledge, a general proof of this implication has not been established in the literature.
From this perspective, $\sigma$–generated calculi may be viewed as isolating the minimal first–order, braid–controlled structure underlying vertical geometry on Hopf–Galois extensions. They capture those aspects of the complete calculus that are essential for the descent of connections and vertical maps, while remaining applicable to arbitrary principal comodule algebras and without imposing higher–order compatibility or exactness conditions.
4 Obstructions to braiding-generated calculi
In the previous section we introduced the notion of $\sigma$-generated first-order differential calculi on principal comodule algebras and proved that, within this class, infinitesimal descent of connection forms, vertical maps hold. The $\sigma$-generation condition is therefore sufficient for descent, but it is by no means automatic.
In this section we identify several necessary conditions for the existence of $\sigma$-generated calculi. Equivalently, we describe obstructions that rule out $\sigma$-generation without requiring explicit computation of the Đurđević braiding. All conditions are expressed purely in terms of the Hopf–Galois data and the balanced calculus.
We first observe that $\sigma$-generation forbids defining relations that are independent of the vertical data determined by a strong connection.
Proposition 4.1.
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let
$\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A}$
be a $\sigma$-generated first-order differential calculus. Then every balanced defining relation in $N_{A}^{\mathrm{bal}}$ is generated from vertical elements via the action of the Đurđević braiding.
Proof.
By $\sigma$-generation, there exists a strong connection $1$-form $\omega$ and a right ideal $I_{H}\subset H^{+}$ such that
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}\;\subset\;A\otimes_{B}A,$
where the $\sigma$-closure is taken with respect to the Đurđević braiding on $A\otimes_{B}A$.
Hence every element of $N_{A}^{\mathrm{bal}}$ is obtained from the vertical data $\pi(\omega(I_{H}))$ by repeated application of $\sigma$ and the $A$-bimodule operations. In particular, no balanced defining relation that is independent of the vertical data can belong to $N_{A}^{\mathrm{bal}}$. ∎
The Đurđević braiding is constructed directly from the translation map
$\tau:H\longrightarrow A\otimes_{B}A,\qquad\mathrm{can}(\tau(h))=1\otimes h.$
It is therefore natural that $\sigma$-generation imposes restrictions on how the image of $\tau$ interacts with the balanced defining relations of a calculus.
Proposition 4.2.
Let $\Omega^{1}(A)$ be $\sigma$-generated with respect to a right ideal $I_{H}\subset H^{+}$. Then
$\langle\tau(I_{H})\rangle_{\sigma}=N_{A}^{\mathrm{bal}}.$
Proof.
Since the calculus is $\sigma$-generated, there exists a strong connection $1$-form $\omega$ such that
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}\;\subset\;A\otimes_{B}A.$
For $h^{+}\in I_{H}\subset H^{+}$,
$\pi(\omega(h^{+}))=\tau(h^{+}).$
Therefore $\pi(\omega(I_{H}))=\tau(I_{H})$, and taking $\sigma$-closures yields
$\langle\tau(I_{H})\rangle_{\sigma}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}=N_{A}^{\mathrm{bal}}.$
Remark 4.3.
Unlike the bicovariant setting of Woronowicz [15], where first–order differential calculi on a Hopf algebra $H$ are classified by right adjoint coaction invariant ideal $I\subset H^{+}$, $\sigma$–generation does not, in general, force the defining subspace $I_{H}\subset H^{+}$ to be a right adjoint coaction invariant ideal of $H$. This reflects the fact that $\sigma$–generated calculi form a strictly more general class than bicovariant differential calculi, as they require only right HH–covariance and compatibility with the Đurđević braiding rather than full bicovariance.
5 The quantum Hopf fibration over the standard Podleś sphere
In this section we treat the standard Podleś sphere as an explicit example of a principal comodule algebra and apply the general existence theorem for $\sigma$–generated differential calculus. Unlike the general quantum flag manifold case, principality can here be proved constructively using the $\mathbb{Z}$–grading induced by the U(1)U(1)–coaction.
The U(1)U(1)–coaction and $\mathbb{Z}$–grading
Let
$A:=\mathcal{O}_{q}(SU(2)),\qquad H:=\mathcal{O}(U(1)),$
where $H$ is generated by a unitary group-like element $t$. The standard right HH–coaction on $A$ is given on generators by
$\delta(\alpha)=\alpha\otimes t,\qquad\delta(\gamma)=\gamma\otimes t,$
$\delta(\alpha^{*})=\alpha^{*}\otimes t^{-1},\qquad\delta(\gamma^{*})=\gamma^{*}\otimes t^{-1}.$
This coaction induces a $\mathbb{Z}$–grading
$A=\bigoplus_{n\in\mathbb{Z}}A_{n},\qquad\delta(a)=a\otimes t^{n}\ \text{for }a\in A_{n},$
whose degree–zero subalgebra is
$B:=A^{\mathrm{co}(H)}=\mathcal{O}_{q}(S^{2}),$
the standard Podleś sphere.
It is well known, going back to Podleś [14], that this grading is strong and that
$A=\bigoplus_{n\in\mathbb{Z}}B\,u_{n},$
where $u_{n}=\alpha^{n}$ for $n\geq 0$ and $u_{-n}=(\alpha^{*})^{n}$ for $n>0$. Every element $a\in A$ admits a unique finite decomposition
$a=\sum_{n\in\mathbb{Z}}b_{n}\,u_{n},\qquad b_{n}\in B.$
Proposition 5.1.
The inclusion
$\mathcal{O}_{q}(S^{2})\subset\mathcal{O}_{q}(SU(2))$
defines a principal right $\mathcal{O}(U(1))$–comodule algebra in the sense of [6].
Proof.
Define a linear map
$s:A\longrightarrow B\otimes A,\qquad s\!\left(\sum_{n}b_{n}\,u_{n}\right)=\sum_{n}b_{n}\otimes u_{n}.$
By construction, $s$ is left BB–linear. Moreover, since $\delta(u_{n})=u_{n}\otimes t^{n}$, one computes
$(\mathrm{id}\otimes\delta)\circ s=(s\otimes\mathrm{id})\circ\delta,$
so that $s$ is right HH–colinear. Finally, letting $\mu:B\otimes A\to A$ denote multiplication, one has
$\mu\circ s(a)=a\quad\text{for all }a\in A.$
Thus $s$ is a left BB–linear, right HH–colinear splitting of the multiplication map.
Since the $\mathbb{Z}$–grading on $A$ is strong, the canonical map
$\mathrm{can}:A\otimes_{B}A\longrightarrow A\otimes H$
is bijective. Therefore all axioms of a principal right HH–comodule algebra in the sense of [6] are satisfied. ∎
We now apply the general existence result for $\sigma$–generated calculi proved earlier in this paper.
Proposition 5.2.
Let $A=\mathcal{O}_{q}(SU(2))$ and $H=\mathcal{O}(U(1))$. For any right ideal $I_{H}\subset H^{+}$, there exists a right HH–covariant first-order differential calculus
$\Omega^{1}(A)=\Omega^{1}_{\mathrm{u}}(A)/N_{A}$
on $A$ which is $\sigma$–generated with respect to $I_{H}$, i.e.
$N_{A}^{\mathrm{bal}}=\big\langle\pi(\omega(I_{H}))\big\rangle_{\sigma}\;\subset\;A\otimes_{B}A,\qquad B=\mathcal{O}_{q}(S^{2}),$
where $\sigma$ denotes the Đurđević braiding associated to the Hopf–Galois extension $B\subset A$.
Proof.
By Proposition 5.1, AA is a principal right HH–comodule algebra. Hence AA admits a strong connection and an associated connection 11–form $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)$. The statement now follows directly from the general existence theorem for $\sigma$–generated calculi established in Proposition 3.7. ∎
Remark 5.3.
As noted above, the explicit splitting map for the Podleś sphere relies on the strong $\mathbb{Z}$–grading of $\mathcal{O}_{q}(SU(2))$ induced by the $\mathcal{O}(U(1))$–coaction. For higher–rank quantum flag manifolds, such as $\mathcal{O}_{q}(\mathbb{CP}^{n})$ with $n\geq 2$, the corresponding coaction still induces a natural $\mathbb{Z}$–grading on $\mathcal{O}_{q}(SU(n+1))$, but this grading is no longer strong: the homogeneous components are nontrivial finitely generated projective modules over the base algebra. As a consequence, no direct algebraic construction of a splitting map analogous to the Podleś case is known, and principality is established by abstract Hopf–Galois methods.
Acknowledgements
The author would like to thank Thomas Weber for suggesting the reference for Đurđević braiding.